# Common Trends and Long-Run Multipliers in Nonlinear Structural VARs

### v1

@@ Appendix A Auxiliary lemmas
    $f(x)-f(x^{\prime})=\beta^{\mathsf{T}}(x-x^{\prime})$ (A.1)
@@
Recall the definition of a bi-Lipschitz function given in 2.3.
Lemma A.4.
@@
Appendix B Proofs of auxiliary lemmas
@@
$\boldsymbol{\chi}(\boldsymbol{z})\coloneqq\begin{bmatrix}\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}\boldsymbol{z}\\ \boldsymbol{\theta}(z)+\boldsymbol{D}_{0}\boldsymbol{z}\end{bmatrix}=\begin{bmatrix}\alpha_{\perp}^{\mathsf{T}}[z-E^{\mathsf{T}}\boldsymbol{\zeta}]\\ \theta(z)\\ \boldsymbol{g}(z)+D\boldsymbol{\zeta}\end{bmatrix}$ (B.1)
where $\boldsymbol{z}=(z^{\mathsf{T}},\boldsymbol{\zeta}^{\mathsf{T}})^{\mathsf{T}}$ for $z\in\mathbb{R}^{p}$ and $\boldsymbol{\zeta}\in\mathbb{R}^{p(k-1)}$. By evaluating (LABEL:bold-chi) when $\boldsymbol{\zeta}=0$, and applying A.1, we see that (iii) implies that $\theta$ and $\boldsymbol{g}$ are Lipschitz, with Lipschitz constant depending only on $\overline{b}$. Hence $g_{i}$ is Lipschitz for $i\in\{1,\ldots,k-1\}$, and moreover so too is
@@
$\boldsymbol{\chi}(\boldsymbol{z}_{0})-\boldsymbol{\chi}(\boldsymbol{z}_{1})=\begin{bmatrix}\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}\\ \boldsymbol{\beta}^{\mathsf{T}}\end{bmatrix}(\boldsymbol{z}_{0}-\boldsymbol{z}_{1}).$ (B.2)
@@
$\lVert\chi(z_{0})-\chi(z_{1})\rVert$ $=\lVert\boldsymbol{\chi}(\boldsymbol{z}_{0})-\boldsymbol{\chi}(\boldsymbol{z}_{1})\rVert\leq C_{0}\lVert\boldsymbol{z}_{0}-\boldsymbol{z}_{1}\rVert.$ (B.3)
@@
$\lVert\boldsymbol{z}_{0}-\boldsymbol{z}_{1}\rVert=\left\|\begin{bmatrix}z_{0}-z_{1}\\ -D^{-1}[\boldsymbol{g}(z_{0})-\boldsymbol{g}(z_{1})]\end{bmatrix}\right\|\leq C_{1}\lVert z_{0}-z_{1}\rVert.$ (B.4)
By A.3, the matrix on the r.h.s. of (LABEL:boldco) is invertible; hence
@@
$\geq\inf_{\boldsymbol{\beta}\in{\cal B}}\left\|\begin{bmatrix}\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}\\ \boldsymbol{\beta}^{\mathsf{T}}\end{bmatrix}^{-1}\right\|^{-1}\lVert\boldsymbol{z}_{0}-\boldsymbol{z}_{1}\rVert$ (B.7)
$=\left(\sup_{\boldsymbol{\beta}\in{\cal B}}\left\|\begin{bmatrix}\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}\\ \boldsymbol{\beta}^{\mathsf{T}}\end{bmatrix}^{-1}\right\|\right)^{-1}\lVert\boldsymbol{z}_{0}-\boldsymbol{z}_{1}\rVert$ (B.10)
@@
Finally, suppose that $f_{0}$ is not the identity, but merely a homeomorphism (as per (i)). By A.2, the model $(c^{\ast},\{f_{i}^{\ast}\}_{i=0}^{k})$, with $c^{\ast}=c$ and $f_{i}^{\ast}=f_{i}\circ f_{0}^{-1}$ for $i\in\{0,\ldots,k\}$ is in class ${\cal M}_{r}^{\ast}(\alpha,\overline{b},\overline{\rho})$, and it follows by the preceding that each $\{f_{i}^{\ast}\}_{i=1}^{k}$ is Lipschitz, and the associated $\chi^{\ast}$ is bi-Lipschitz. Clearly $f_{i}=f_{i}^{\ast}\circ f_{0}$ is continuous for each $i\in\{1,\ldots,k\}$. Moreover $\chi^{\ast}=\chi\circ f_{0}^{-1}$, whence $\chi=\chi^{\ast}\circ f_{0}$ is also a homeomorphism. If $f_{0}$ is bi-Lipschitz with bi-Lipschitz constant $C_{f_{0}}$, then $\chi=\chi^{\ast}\circ f_{0}$ is bi-Lipschitz, with bi-Lipschitz constant depending only on $C_{f_{0}}$, $\alpha$, $\overline{b}$ and $\overline{\rho}$. ∎
Appendix C Proofs of high-level results
@@
We note that $\chi$ is a homeomorphism by A.4.
@@
$\Delta\boldsymbol{z}_{t}$ $=\boldsymbol{c}+\boldsymbol{\pi}(z_{t-1})+\boldsymbol{D}\boldsymbol{z}_{t-1}+\boldsymbol{u}_{t}$ (C.1)
@@
$\boldsymbol{\pi}(z)=\begin{bmatrix}\pi(z)+g_{1}(z)\\ \boldsymbol{g}(z)\end{bmatrix}=\begin{bmatrix}\alpha&E^{\mathsf{T}}\\ 0&I_{p(k-1)}\end{bmatrix}\begin{bmatrix}\theta(z)\\ \boldsymbol{g}(z)\end{bmatrix}=\boldsymbol{\alpha}\boldsymbol{\theta}(z),$ (C.2)
@@
$\boldsymbol{c}=\begin{bmatrix}c\\ 0_{p(k-1)}\end{bmatrix}=\begin{bmatrix}\alpha&E^{\mathsf{T}}\\ 0&I_{p(k-1)}\end{bmatrix}\begin{bmatrix}\mu\\ 0_{p(k-1)}\end{bmatrix}\eqqcolon\boldsymbol{\alpha}\boldsymbol{\mu}.$ (C.3)
@@
$\boldsymbol{\xi}_{t}=(I_{p(k-1)+r}+\boldsymbol{\beta}_{t}^{\mathsf{T}}\boldsymbol{\alpha})\boldsymbol{\xi}_{t-1}+\boldsymbol{\beta}_{t}^{\mathsf{T}}\boldsymbol{u}_{t}$ (C.4)
@@
$\boldsymbol{\xi}_{t}=\boldsymbol{\mu}+\boldsymbol{\theta}(z_{t-1})+\boldsymbol{D}_{0}\boldsymbol{z}_{t-1}=\begin{bmatrix}\mu+\theta(z_{t})\\ \boldsymbol{g}(z_{t})+D\boldsymbol{\zeta}_{t-1}\end{bmatrix}=\begin{bmatrix}\xi_{t}\\ \Delta\boldsymbol{\zeta_{t}}\end{bmatrix}$ (C.5)
@@
$\theta(z_{t})=-\mu+S_{\theta}^{\mathsf{T}}\boldsymbol{\xi}_{t}$ (C.6)
@@
$=\alpha_{\perp}^{\mathsf{T}}\overline{h}(\vec{z}_{0})+\alpha_{\perp}^{\mathsf{T}}\sum_{s=1}^{t}u_{s}.$ (C.7)
@@
$-D^{-1}=\begin{bmatrix}I_{p}&-I_{p}\\ &\ddots&\ddots\\ &&I_{p}&-I_{p}\\ &&&I_{p}\end{bmatrix}^{-1}=\begin{bmatrix}I_{p}&I_{p}&I_{p}&I_{p}\\ &\ddots&\vdots&\vdots\\ &&I_{p}&I_{p}\\ &&&I_{p}\end{bmatrix}.$ (C.8)
@@
$\psi(z_{t})=\alpha_{\perp}^{\mathsf{T}}\overline{h}(\vec{z}_{0})+\alpha_{\perp}^{\mathsf{T}}\sum_{s=1}^{t}u_{s}+\alpha_{\perp}^{\mathsf{T}}S_{\psi}^{\mathsf{T}}\boldsymbol{\xi}_{t}$ (C.9)
@@
i.e. it is generated by a model $(c^{\ast},\{f_{i}^{\ast}\}_{i=0}^{k})$ with $f_{0}^{\ast}=I_{p}$. By A.2, this model belongs to class ${\cal M}_{r}^{\ast}$, with the same $\alpha$ and $\mu$. Therefore, by the arguments given above, (LABEL:gjrt-z) holds in the form
@@
It additionally follows from A.2 that the l.h.s. is equal to $\chi(z_{t})$, while
@@
$\boldsymbol{\xi}_{t}^{\ast}=\begin{bmatrix}\mu+\theta^{\ast}(z_{t}^{\ast})\\ \Delta\boldsymbol{\zeta}_{t}^{\ast}\end{bmatrix}=\begin{bmatrix}\mu+\theta(z_{t})\\ \Delta\boldsymbol{\zeta}_{t}\end{bmatrix}=\boldsymbol{\xi}_{t}$ (C.10)
@@
where $\boldsymbol{\beta}_{t}\in{\cal B}^{\ast}$ for each $t\in\mathbb{N}$, it then follows by (LABEL:xi-ast-xi) and A.2 that
@@
Without loss of generality, we may take $\tau=1$. We note that $\chi$ is a homeomorphism by A.4.
@@
Appendix D Proofs relating to examples
D.1 Linear systems
@@
$I_{pk}+\boldsymbol{\alpha}\boldsymbol{\beta}^{\mathsf{T}}=I_{pk}+\begin{bmatrix}\alpha&E^{\mathsf{T}}\\ 0&I_{p(k-1)}\end{bmatrix}\begin{bmatrix}\beta^{\mathsf{T}}\Phi_{0}^{-1}&0\\ \boldsymbol{\Gamma}\Phi_{0}^{-1}&D\end{bmatrix}=\begin{bmatrix}I_{p}+(\alpha\beta^{\mathsf{T}}+\Gamma_{1})\Phi_{0}^{-1}&E^{\mathsf{T}}D\\ \boldsymbol{\Gamma}\Phi_{0}^{-1}&I_{p(k-1)}+D\end{bmatrix}$ (D.1)
@@
It therefore follows from A.4 that the matrix on the r.h.s. is invertible. Since
@@
D.2 Piecewise affine systems
Recall from 7 that $f:\mathbb{R}^{p}\rightarrow\mathbb{R}^{p}$ is piecewise affine if it is a continuous function of the form
$f(x)=\sum_{\ell=1}^{L}\mathbf{1}\{x\in\mathscr{X}^{(\ell)}\}(\phi^{(\ell)}+\Phi^{(\ell)}x),$ (D.2)
@@ D.2 Piecewise affine systems
    $\mathscr{Z}^{(\ell)}=\{z\in\mathbb{R}^{p}\mid a^{\mathsf{T}}z\in(\tau_{\ell-1},\tau_{\ell}]\}.$ (D.3)
@@
$=[\phi(x^{\prime\prime})-\phi(x^{\prime})]+\Phi(x^{\prime})(x^{\prime\prime}-x^{\prime})+[\Phi(x^{\prime\prime})-\Phi(x^{\prime})]x^{\prime\prime}.$ (D.4)
@@
$0=\lim_{\delta\downarrow\delta_{i}}f[x(\delta)]-\lim_{\delta\uparrow\delta_{i}}f[x(\delta)]=[\phi(x_{i})-\phi(x_{i-1})]+[\Phi(x_{i})-\Phi(x_{i-1})]x(\delta_{i}).$ (D.5)
@@
$x^{\prime\prime}-x(\delta_{i})=x^{\prime\prime}-[(1-\delta_{i})x^{\prime}+\delta_{i}x^{\prime\prime}]=(1-\delta_{i})(x^{\prime\prime}-x^{\prime}),$ (D.6)
@@
$=\frac{1}{1+a^{\mathsf{T}}(\Phi^{(1)})^{-1}n^{(\ell)}}a^{\mathsf{T}}(\Phi^{(1)})^{-1}=c_{\ell}a^{\mathsf{T}}(\Phi^{(1)})^{-1}.$ (D.10)
@@
$\delta_{n}\coloneqq\inf_{\lVert z\rVert\geq n}\lVert f(z)\rVert\rightarrow\infty$ (D.11)
@@
$\pi(z)=\sum_{\ell=1}^{L}\mathbf{1}^{(\ell)}(z)(\bar{\pi}^{(\ell)}+\Pi^{(\ell)}z)=\alpha\sum_{\ell=1}^{L}\mathbf{1}^{(\ell)}(z)(\bar{\mu}^{(\ell)}+\beta^{(\ell)\mathsf{T}}z)=\alpha\theta(z)$ (D.12)
@@
$=\sum_{\ell=1}^{L}\mathbf{1}^{(\ell)}(z)[\bar{\gamma}_{j}^{(\ell)}+\Gamma_{j}^{(\ell)}z].$ (D.13)
@@
$\boldsymbol{\theta}(z)=\begin{bmatrix}\theta(z)\\ \boldsymbol{g}(z)\end{bmatrix}=\sum_{\ell=1}^{L}\mathbf{1}^{(\ell)}(z)\left(\begin{bmatrix}\bar{\mu}^{(\ell)}\\ \bar{\boldsymbol{\gamma}}^{(\ell)}\end{bmatrix}+\begin{bmatrix}\beta^{(\ell)\mathsf{T}}\\ \boldsymbol{\Gamma}^{(\ell)}\end{bmatrix}z\right)\eqqcolon m(z)+M(z)z,$ (D.14)
@@
it follows from D.2(ii) that
@@
is piecewise affine on $\{\mathscr{Y}^{(\ell)}\times\mathbb{R}^{p(k-1)}\}_{\ell=1}^{L}$, which provide a convex partition of $\mathbb{R}^{kp}$. Therefore by D.2, for any $\boldsymbol{y},\boldsymbol{y}^{\prime}\in\mathbb{R}^{kp}$,
@@
(i). In view of (i), D.3 yields that $f_{0}$ is a homeomorphism. (Note that the subsequent arguments only require that $(c,\{f_{i}\}_{i=0}^{k})$ be piecewise affine, so that the conclusions of the theorem also hold when it is directly assumed that $f_{0}$ is a homeomorphism.)
@@
(iii). 7.1 yields that (LABEL:gradient) is satisfied for ${\cal B}\subset\mathbb{R}^{p(k-1)+r}$ such that
@@
Finally, the claimed form for $\chi$ follows by noting that $\theta(z)$ must be as in (LABEL:pi-pwa) in the proof of 7.1, while (LABEL:ct-def) here implies that
@@
we we have used the expression for $g_{i}$ given in (LABEL:ga-pwa) in the proof of 7.1. ∎
Proof of 7.2.
@@
where $\boldsymbol{\beta}_{u}\in{\cal B}$ for each $u\in\mathbb{R}^{p}$, for ${\cal B}$ the set of gradients associated to $(c,\{f_{i}\}_{i=0}^{k})$. Since ${\cal B}$ is convex by 7.1, $\overline{\boldsymbol{\beta}}\in{\cal B}$. Hence (iii) holds with ${\cal B}_{K}={\cal B}$. ∎

### v2

@@ Appendix B Auxiliary lemmas
    $f(x)-f(x^{\prime})=\beta^{\mathsf{T}}(x-x^{\prime})$ (B.1)
@@
Recall the definition of a bi-Lipschitz function given in 2.2.
Lemma B.4.
@@
Appendix C Proofs of auxiliary lemmas
@@
$\boldsymbol{\chi}(\boldsymbol{z})\coloneqq\begin{bmatrix}\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}\boldsymbol{z}\\ \boldsymbol{\theta}(z)+\boldsymbol{D}_{0}\boldsymbol{z}\end{bmatrix}=\begin{bmatrix}\alpha_{\perp}^{\mathsf{T}}[z-E^{\mathsf{T}}\boldsymbol{\zeta}]\\ \theta(z)\\ \boldsymbol{g}(z)+D\boldsymbol{\zeta}\end{bmatrix}$ (C.1)
where $\boldsymbol{z}=(z^{\mathsf{T}},\boldsymbol{\zeta}^{\mathsf{T}})^{\mathsf{T}}$ for $z\in\mathbb{R}^{p}$ and $\boldsymbol{\zeta}\in\mathbb{R}^{p(k-1)}$. By evaluating (LABEL:bold-chi) when $\boldsymbol{\zeta}=0$, and applying B.1, we see that (iii) implies that $\theta$ and $\boldsymbol{g}$ are Lipschitz, with Lipschitz constant depending only on $\overline{b}$. Hence $g_{i}$ is Lipschitz for $i\in\{1,\ldots,k-1\}$, and moreover so too is
@@
$\boldsymbol{\chi}(\boldsymbol{z}_{0})-\boldsymbol{\chi}(\boldsymbol{z}_{1})=\begin{bmatrix}\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}\\ \boldsymbol{\beta}^{\mathsf{T}}\end{bmatrix}(\boldsymbol{z}_{0}-\boldsymbol{z}_{1}).$ (C.2)
@@
$\lVert\chi(z_{0})-\chi(z_{1})\rVert$ $=\lVert\boldsymbol{\chi}(\boldsymbol{z}_{0})-\boldsymbol{\chi}(\boldsymbol{z}_{1})\rVert\leq C_{0}\lVert\boldsymbol{z}_{0}-\boldsymbol{z}_{1}\rVert.$ (C.3)
@@
$\lVert\boldsymbol{z}_{0}-\boldsymbol{z}_{1}\rVert=\left\|\begin{bmatrix}z_{0}-z_{1}\\ -D^{-1}[\boldsymbol{g}(z_{0})-\boldsymbol{g}(z_{1})]\end{bmatrix}\right\|\leq C_{1}\lVert z_{0}-z_{1}\rVert.$ (C.4)
By B.3, the matrix on the r.h.s. of (LABEL:boldco) is invertible; hence
@@
$\geq\inf_{\boldsymbol{\beta}\in{\cal B}}\left\|\begin{bmatrix}\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}\\ \boldsymbol{\beta}^{\mathsf{T}}\end{bmatrix}^{-1}\right\|^{-1}\lVert\boldsymbol{z}_{0}-\boldsymbol{z}_{1}\rVert$
$=\left(\sup_{\boldsymbol{\beta}\in{\cal B}}\left\|\begin{bmatrix}\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}\\ \boldsymbol{\beta}^{\mathsf{T}}\end{bmatrix}^{-1}\right\|\right)^{-1}\lVert\boldsymbol{z}_{0}-\boldsymbol{z}_{1}\rVert$
@@
Finally, suppose that $f_{0}$ is not the identity, but merely a homeomorphism (as per (i)). By B.2, the model $(c^{\ast},\{f_{i}^{\ast}\}_{i=0}^{k})$, with $c^{\ast}=c$ and $f_{i}^{\ast}=f_{i}\circ f_{0}^{-1}$ for $i\in\{0,\ldots,k\}$ is in class ${\cal M}_{r}^{\ast}(\alpha,\overline{b},\overline{\rho})$, and it follows by the preceding that each $\{f_{i}^{\ast}\}_{i=1}^{k}$ is Lipschitz, and the associated $\chi^{\ast}$ is bi-Lipschitz. Clearly $f_{i}=f_{i}^{\ast}\circ f_{0}$ is continuous for each $i\in\{1,\ldots,k\}$. Moreover $\chi^{\ast}=\chi\circ f_{0}^{-1}$, whence $\chi=\chi^{\ast}\circ f_{0}$ is also a homeomorphism. If $f_{0}$ is bi-Lipschitz with bi-Lipschitz constant $C_{f_{0}}$, then $\chi=\chi^{\ast}\circ f_{0}$ is bi-Lipschitz, with bi-Lipschitz constant depending only on $C_{f_{0}}$, $\alpha$, $\overline{b}$ and $\overline{\rho}$. ∎
Appendix D Proofs of high-level results
@@
We note that $\chi$ is a homeomorphism by B.4.
@@
$\Delta\boldsymbol{z}_{t}$ $=\boldsymbol{c}+\boldsymbol{\pi}(z_{t-1})+\boldsymbol{D}\boldsymbol{z}_{t-1}+\boldsymbol{u}_{t}$ (D.1)
@@
$\boldsymbol{\pi}(z)=\begin{bmatrix}\pi(z)+g_{1}(z)\\ \boldsymbol{g}(z)\end{bmatrix}=\begin{bmatrix}\alpha&E^{\mathsf{T}}\\ 0&I_{p(k-1)}\end{bmatrix}\begin{bmatrix}\theta(z)\\ \boldsymbol{g}(z)\end{bmatrix}=\boldsymbol{\alpha}\boldsymbol{\theta}(z),$ (D.2)
@@
$\boldsymbol{c}=\begin{bmatrix}c\\ 0_{p(k-1)}\end{bmatrix}=\begin{bmatrix}\alpha&E^{\mathsf{T}}\\ 0&I_{p(k-1)}\end{bmatrix}\begin{bmatrix}\mu\\ 0_{p(k-1)}\end{bmatrix}\eqqcolon\boldsymbol{\alpha}\boldsymbol{\mu}.$ (D.3)
@@
$\boldsymbol{\xi}_{t}=(I_{p(k-1)+r}+\boldsymbol{\beta}_{t}^{\mathsf{T}}\boldsymbol{\alpha})\boldsymbol{\xi}_{t-1}+\boldsymbol{\beta}_{t}^{\mathsf{T}}\boldsymbol{u}_{t}$ (D.4)
@@
$\boldsymbol{\xi}_{t}=\boldsymbol{\mu}+\boldsymbol{\theta}(z_{t-1})+\boldsymbol{D}_{0}\boldsymbol{z}_{t-1}=\begin{bmatrix}\mu+\theta(z_{t})\\ \boldsymbol{g}(z_{t})+D\boldsymbol{\zeta}_{t-1}\end{bmatrix}=\begin{bmatrix}\xi_{t}\\ \Delta\boldsymbol{\zeta_{t}}\end{bmatrix}$ (D.5)
@@
$\theta(z_{t})=-\mu+S_{\theta}^{\mathsf{T}}\boldsymbol{\xi}_{t}$ (D.6)
@@
$=\alpha_{\perp}^{\mathsf{T}}\overline{h}(\vec{z}_{0})+\alpha_{\perp}^{\mathsf{T}}\sum_{s=1}^{t}u_{s}.$ (D.7)
@@
$-D^{-1}=\begin{bmatrix}I_{p}&-I_{p}\\ &\ddots&\ddots\\ &&I_{p}&-I_{p}\\ &&&I_{p}\end{bmatrix}^{-1}=\begin{bmatrix}I_{p}&I_{p}&I_{p}&I_{p}\\ &\ddots&\vdots&\vdots\\ &&I_{p}&I_{p}\\ &&&I_{p}\end{bmatrix}.$ (D.8)
@@
$\psi(z_{t})=\alpha_{\perp}^{\mathsf{T}}\overline{h}(\vec{z}_{0})+\alpha_{\perp}^{\mathsf{T}}\sum_{s=1}^{t}u_{s}+\alpha_{\perp}^{\mathsf{T}}S_{\psi}^{\mathsf{T}}\boldsymbol{\xi}_{t}$ (D.9)
@@
i.e. it is generated by a model $(c^{\ast},\{f_{i}^{\ast}\}_{i=0}^{k})$ with $f_{0}^{\ast}=I_{p}$. By B.2, this model belongs to class ${\cal M}_{r}^{\ast}$, with the same $\alpha$ and $\mu$. Therefore, by the arguments given above, (LABEL:gjrt-z) holds in the form
@@
It additionally follows from B.2 that the l.h.s. is equal to $\chi(z_{t})$, while
@@
$\boldsymbol{\xi}_{t}^{\ast}=\begin{bmatrix}\mu+\theta^{\ast}(z_{t}^{\ast})\\ \Delta\boldsymbol{\zeta}_{t}^{\ast}\end{bmatrix}=\begin{bmatrix}\mu+\theta(z_{t})\\ \Delta\boldsymbol{\zeta}_{t}\end{bmatrix}=\boldsymbol{\xi}_{t}$ (D.10)
@@
where $\boldsymbol{\beta}_{t}\in{\cal B}^{\ast}$ for each $t\in\mathbb{N}$, it then follows by (LABEL:xi-ast-xi) and B.2 that
@@
Without loss of generality, we may take $\tau=1$. We note that $\chi$ is a homeomorphism by B.4.
@@
Appendix E Proofs relating to examples
E.1 Linear systems
@@
$I_{pk}+\boldsymbol{\alpha}\boldsymbol{\beta}^{\mathsf{T}}=I_{pk}+\begin{bmatrix}\alpha&E^{\mathsf{T}}\\ 0&I_{p(k-1)}\end{bmatrix}\begin{bmatrix}\beta^{\mathsf{T}}\Phi_{0}^{-1}&0\\ \boldsymbol{\Gamma}\Phi_{0}^{-1}&D\end{bmatrix}=\begin{bmatrix}I_{p}+(\alpha\beta^{\mathsf{T}}+\Gamma_{1})\Phi_{0}^{-1}&E^{\mathsf{T}}D\\ \boldsymbol{\Gamma}\Phi_{0}^{-1}&I_{p(k-1)}+D\end{bmatrix}$ (E.1)
@@
It therefore follows from B.4 that the matrix on the r.h.s. is invertible. Since
@@
E.2 Piecewise affine systems
Recall from 6 that $f:\mathbb{R}^{p}\rightarrow\mathbb{R}^{p}$ is piecewise affine if it is a continuous function of the form
$f(x)=\sum_{\ell=1}^{L}\mathbf{1}\{x\in\mathscr{X}^{(\ell)}\}(\phi^{(\ell)}+\Phi^{(\ell)}x),$ (E.2)
@@ E.2 Piecewise affine systems
    $\mathscr{Z}^{(\ell)}=\{z\in\mathbb{R}^{p}\mid a^{\mathsf{T}}z\in(\tau_{\ell-1},\tau_{\ell}]\}.$ (E.3)
@@
$=[\phi(x^{\prime\prime})-\phi(x^{\prime})]+\Phi(x^{\prime})(x^{\prime\prime}-x^{\prime})+[\Phi(x^{\prime\prime})-\Phi(x^{\prime})]x^{\prime\prime}.$ (E.4)
@@
$0=\lim_{\delta\downarrow\delta_{i}}f[x(\delta)]-\lim_{\delta\uparrow\delta_{i}}f[x(\delta)]=[\phi(x_{i})-\phi(x_{i-1})]+[\Phi(x_{i})-\Phi(x_{i-1})]x(\delta_{i}).$ (E.5)
@@
$x^{\prime\prime}-x(\delta_{i})=x^{\prime\prime}-[(1-\delta_{i})x^{\prime}+\delta_{i}x^{\prime\prime}]=(1-\delta_{i})(x^{\prime\prime}-x^{\prime}),$ (E.6)
@@
$=\frac{1}{1+a^{\mathsf{T}}(\Phi^{(1)})^{-1}n^{(\ell)}}a^{\mathsf{T}}(\Phi^{(1)})^{-1}=c_{\ell}a^{\mathsf{T}}(\Phi^{(1)})^{-1}.$ (E.10)
@@
$\delta_{n}\coloneqq\inf_{\lVert z\rVert\geq n}\lVert f(z)\rVert\rightarrow\infty$ (E.11)
@@
$\pi(z)=\sum_{\ell=1}^{L}\mathbf{1}^{(\ell)}(z)(\bar{\pi}^{(\ell)}+\Pi^{(\ell)}z)=\alpha\sum_{\ell=1}^{L}\mathbf{1}^{(\ell)}(z)(\bar{\mu}^{(\ell)}+\beta^{(\ell)\mathsf{T}}z)=\alpha\theta(z)$ (E.12)
@@
$=\sum_{\ell=1}^{L}\mathbf{1}^{(\ell)}(z)[\bar{\gamma}_{j}^{(\ell)}+\Gamma_{j}^{(\ell)}z].$ (E.13)
@@
$\boldsymbol{\theta}(z)=\begin{bmatrix}\theta(z)\\ \boldsymbol{g}(z)\end{bmatrix}=\sum_{\ell=1}^{L}\mathbf{1}^{(\ell)}(z)\left(\begin{bmatrix}\bar{\mu}^{(\ell)}\\ \bar{\boldsymbol{\gamma}}^{(\ell)}\end{bmatrix}+\begin{bmatrix}\beta^{(\ell)\mathsf{T}}\\ \boldsymbol{\Gamma}^{(\ell)}\end{bmatrix}z\right)\eqqcolon m(z)+M(z)z,$ (E.14)
@@
it follows from E.2(ii) that
@@
is piecewise affine on $\{\mathscr{Y}^{(\ell)}\times\mathbb{R}^{p(k-1)}\}_{\ell=1}^{L}$, which provide a convex partition of $\mathbb{R}^{kp}$. Therefore by E.2, for any $\boldsymbol{y},\boldsymbol{y}^{\prime}\in\mathbb{R}^{kp}$,
@@
(i). In view of (i), E.3 yields that $f_{0}$ is a homeomorphism. (Note that the subsequent arguments only require that $(c,\{f_{i}\}_{i=0}^{k})$ be piecewise affine, so that the conclusions of the theorem also hold when it is directly assumed that $f_{0}$ is a homeomorphism.)
@@
(iii). 6.1 yields that (LABEL:gradient) is satisfied for ${\cal B}\subset\mathbb{R}^{p(k-1)+r}$ such that
@@
Finally, the claimed form for $\chi$ follows by noting that $\theta(z)$ must be as in (LABEL:pi-pwa) in the proof of 6.1, while (LABEL:ct-def) here implies that
@@
we we have used the expression for $g_{i}$ given in (LABEL:ga-pwa) in the proof of 6.1. ∎
Proof of 6.2.
@@
where $\boldsymbol{\beta}_{u}\in{\cal B}$ for each $u\in\mathbb{R}^{p}$, for ${\cal B}$ the set of gradients associated to $(c,\{f_{i}\}_{i=0}^{k})$. Since ${\cal B}$ is convex by 6.1, $\overline{\boldsymbol{\beta}}\in{\cal B}$. Hence (iii) holds with ${\cal B}_{K}={\cal B}$. ∎